### v3

# The linear targeting problem

Abstract.
For given real or complex $m\times n$ data matrices $X$, $Y$, we investigate when there is a matrix $A$ such that $AX=Y$, and $A$ is invertible, Hermitian, positive (semi)definite, unitary, an orthogonal projection, a reflection, complex symmetric, or normal.
Key words and phrases:
data mapping, source and target matrices, constrained systems of linear equations
2020 Mathematics Subject Classification:
15A06, 15A10, 15A24
SRG partially supported by NSF grant DMS-2054002.
1. Introduction
The linear targeting problem is to construct an $A\in\mathsf{M}_{m}(\mathbb{F})$ such that $AX=Y$, in which $X,Y\in\mathsf{M}_{m\times n}(\mathbb{F})$ are given and $\mathbb{F}=\mathbb{R}$ or $\mathbb{C}$. We say that $X$ is the source, $Y$ is the target, and $A$ is a targeting matrix. If there is an $A\in\mathsf{M}_{m}(\mathbb{F})$ with property $\mathcal{P}$ such that $AX=Y$, we say that $A$ solves the $\mathcal{P}$-targeting problem for $X$ and $Y$. In this paper, we consider the following properties: invertible, Hermitian, positive (semi)definite, unitary, orthogonal projection, reflection, complex symmetric, and normal.
If $AX=Y$, then $\operatorname{null}X\subseteq\operatorname{null}Y$, so this condition (or something that implies it) is found in all of our results. The invertible linear targeting problem is just the row equivalence problem, which has a solution if and only if $\operatorname{null}X=\operatorname{null}Y$ [3, Theorem 4.4.1].
If $m<n$ and $AX=Y$, then after a simultaneous permutation of the columns of $X$ and $Y$ we may assume that $X=[X_{1}\,\,X_{2}]$, $Y=[Y_{1}\,\,Y_{2}]$, $X_{1},Y_{1}\in\mathsf{M}_{m\times k}(\mathbb{F})$, $k\leq m$, and $\operatorname{col}X_{2}\subseteq\operatorname{col}X_{1}$. For example, we could choose $k=\operatorname{rank}X$, but any $k$ between $\operatorname{rank}X$ and $m$ will do. Since there is some $B\in\mathsf{M}_{k\times(n-k)}(\mathbb{F})$ such that $X_{2}=X_{1}B$ [3, Theorem 1.6.21.a], we have $AX_{1}=Y_{1}$, from which it follows that $AX_{2}=A(X_{1}B)=Y_{1}B=Y_{2}$. Thus, in considering the linear targeting problem $AX=Y$ with $X,Y\in\mathsf{M}_{m\times n}(\mathbb{F})$, it suffices to consider only the case $m\geq n\geq 1$.
For any $Z\in\mathsf{M}_{m\times n}(\mathbb{F})$ ($m\times n$ matrices over the field $\mathbb{R}$ or $\mathbb{C}$), $Z^{*}$ is the conjugate transpose of $Z$ and $Z^{\dagger}$ is its Moore–Penrose pseudoinverse; see [3, § 17.4] or [4, 7.3.P7]. The Euclidean norm of $\mathbf{x}\in\mathbb{F}^{n}$ is denoted by $\|\mathbf{x}\|=(\mathbf{x}^{*}\mathbf{x})^{1/2}$. If $A\in\mathsf{M}_{n}(\mathbb{F})$, we write $A\geq 0$ (respectively, $A>0$) if $A\in\mathsf{M}_{n}(\mathbb{F})$ is positive semidefinite (respectively, positive definite), that is, $A$ is Hermitian and $\mathbf{u}^{*}A\mathbf{u}\geq 0$ (respectively, $\mathbf{u}^{*}A\mathbf{u}>0)$ for all nonzero $\mathbf{u}\in\mathbb{F}^{n}$. If $B,S\in\mathsf{M}_{m}(\mathbb{F})$ and $S$ is invertible, then $S^{*}BS$ is congruent to $B$. If $B$ is, respectively, invertible, Hermitian, or positive semidefinite, then $S^{*}BS$ has the same respective properties.
2. Unconstrained linear targeting
A linear targeting problem need not have a solution. For example, $AX=Y$ is impossible if
$X=\begin{bmatrix}1&0\\ 0&0\end{bmatrix}\quad\text{and}\quad Y=\begin{bmatrix}0&1\\ 0&0\end{bmatrix}.$
Since $\operatorname{null}X\subseteq\operatorname{null}(AX)$, it is necessary that $\operatorname{null}X\subseteq\operatorname{null}Y$. A computation with pseudoinverses shows that this condition permits us to identify all solutions of an unconstrained linear targeting problem.
Theorem 2.1.
Let $X,Y\in\mathsf{M}_{m\times n}(\mathbb{F})$ with $m\geq n\geq 1$ and $\operatorname{null}X\subseteq\operatorname{null}Y$.
  1. (a)
    $(YX^{\dagger})X=Y$.
  2. (b)
    If $Z\in\mathsf{M}_{m}(\mathbb{F})$ and $A=YX^{\dagger}+Z(I-XX^{\dagger})$, then $AX=Y$.
  3. (c)
    If $A\in\mathsf{M}_{m}(\mathbb{F})$ and $AX=Y$, then there is a $Z\in\mathsf{M}_{m}(\mathbb{F})$ such that $A=YX^{\dagger}+Z(I-XX^{\dagger})$.
Proof.
(a) The hypothesis $\operatorname{null}X\subseteq\operatorname{null}Y$ is equivalent to the condition $\operatorname{col}Y^{*}\subseteq\operatorname{col}X^{*}$. Since $X^{\dagger}X$ is the orthogonal projection (necessarily Hermitian) onto $\operatorname{col}X^{*}$, we have $X^{\dagger}XY^{*}=Y^{*}$, and hence $Y=YX^{\dagger}X=(YX^{\dagger})X$.
(b) Compute
$AX$ $=(YX^{\dagger}+Z(I-XX^{\dagger}))X$
$=YX^{\dagger}X+Z(X-XX^{\dagger}X)$
$=Y+Z(X-X)=Y.$
(c) If $AX=Y$, then
$A-YX^{\dagger}$ $=(A-YX^{\dagger})(I-XX^{\dagger}+XX^{\dagger})$
$=(A-YX^{\dagger})(I-XX^{\dagger})+(A-YX^{\dagger})XX^{\dagger}$
$=(A-YX^{\dagger})(I-XX^{\dagger})+(AX-(YX^{\dagger})X)X^{\dagger}$
$=(A-YX^{\dagger})(I-XX^{\dagger})+(Y-Y)X^{\dagger}$
$=(A-YX^{\dagger})(I-XX^{\dagger})=Z(I-XX^{\dagger}),$
in which $Z=A-YX^{\dagger}$. ∎
The following theorem re-examines the unconstrained linear targeting problem in the context of a singular value decomposition of the data matrix $X$. Later, we find it convenient to employ a singular value decomposition of $Y$, with analogous notation. If $X=0$, the linear targeting problem has a solution if and only if $Y=0$, so this case is not interesting.
Theorem 2.2.
Let $X,Y\in\mathsf{M}_{m\times n}(\mathbb{F})$ with $m\geq n\geq 1$ and $X\neq 0$. There is an $A\in\mathsf{M}_{m}(\mathbb{F})$ such that $AX=Y$ if and only if $\operatorname{null}X\subseteq\operatorname{null}Y$.
Proof.
($\Rightarrow$) If $\mathbf{u}\in\operatorname{null}X$, then $Y\mathbf{u}=AX\mathbf{u}=A\mathbf{0}=\mathbf{0}$.
($\Leftarrow$) Let $\operatorname{rank}X=r\geq 1$ and let $X=V\Sigma W^{*}$ be a singular value decomposition, in which $V\in\mathsf{M}_{m}(\mathbb{F})$ and $W\in\mathsf{M}_{n}(\mathbb{F})$ are unitary,
$\Sigma=\begin{bmatrix}\Sigma_{r}&0_{r\times(n-r)}\\ 0_{(m-r)\times r}&0_{(m-r)\times(n-r)}\end{bmatrix}\in\mathsf{M}_{m\times n}(\mathbb{R}),$ (1)
$r\geq 1$, and $\Sigma_{r}\in\mathsf{M}_{r}(\mathbb{R})$ is diagonal, invertible, and positive definite. Partition $W=[W_{1}~{}W_{2}]$ and $V=[V_{1}~{}V_{2}]$, in which $W_{1}\in\mathsf{M}_{n\times r}(\mathbb{F})$ and $V_{1}\in\mathsf{M}_{m\times r}(\mathbb{F})$. If $X$ has full rank, the block column of zeros in (1) is absent, $W=W_{1}$, and $W_{2}$ is absent. If $r<n$, the columns of $W_{2}$ are an orthonormal basis for $\operatorname{null}X$ and the columns of $W_{1}$ are an orthonormal basis for $(\operatorname{null}X)^{\perp}=\operatorname{col}X^{*}$. Since $\operatorname{null}X\subseteq\operatorname{null}Y$, we have $YW_{2}=0$.
If there is a $B\in\mathsf{M}_{m}$ such that
$B\Sigma=V^{*}YW,$ (2)
then $YW=VB\Sigma=(VBV^{*})(V\Sigma W^{*})W=(VBV^{*})XW$ and hence
$(VBV^{*})X=Y.$
Thus, $A=VBV^{*}$ solves the unconstrained targeting problem for $X$ and $Y$. We can partition the right side of (2) in several ways:
$V^{*}YW$ $=V^{*}Y[W_{1}~{}W_{2}]=V^{*}[YW_{1}~{}YW_{2}]=[V^{*}YW_{1}~{}0]$ (3)
$=\begin{bmatrix}V_{1}^{*}YW_{1}&0\\ V_{2}^{*}YW_{1}&0\end{bmatrix}=\begin{bmatrix}Z_{1}&0\\ Z_{2}&0\end{bmatrix}=\begin{bmatrix}Z&0\end{bmatrix}\in\mathsf{M}_{m\times n}(\mathbb{F}),$
in which the three matrices $Z=V^{*}YW_{1}\in\mathsf{M}_{m\times r}(\mathbb{F})$, $Z_{1}=V_{1}^{*}YW_{1}\in\mathsf{M}_{r}(\mathbb{F})$, and $Z_{2}=V_{2}^{*}YW_{1}\in\mathsf{M}_{(m-r)\times r}(\mathbb{F})$ are determined by the data $X$ and $Y$. Define $B_{1}=Z\Sigma_{r}^{-1}\in\mathsf{M}_{m\times r}(\mathbb{F})$ and let
$B=[B_{1}\,\,B_{2}]=\begin{bmatrix}Z\Sigma_{r}^{-1}&B_{2}\end{bmatrix}=\begin{bmatrix}Z_{1}\Sigma_{r}^{-1}&R\\ Z_{2}\Sigma_{r}^{-1}&S\end{bmatrix},$ (4)
in which $R\in\mathsf{M}_{r\times(n-r)}(\mathbb{F})$ and $S\in\mathsf{M}_{n-r}(\mathbb{F})$ are arbitrary. Then
$B\Sigma=\begin{bmatrix}Z\Sigma_{r}^{-1}&B_{2}\end{bmatrix}\begin{bmatrix}\Sigma_{r}&0\\ 0&0\end{bmatrix}=\begin{bmatrix}Z&0\end{bmatrix}=V^{*}YW,$
so (2) is satisfied for any choice of $B_{2}$. ∎
If a property $\mathcal{P}$ is invariant under unitary similarity, the preceding theorem shows that the $\mathcal{P}$-targeting problem is equivalent to a matrix completion problem. The source and target matrices determine $B_{1}$; if we can choose $B_{2}$ so that $B$ (and hence also $VBV^{*})$ has property $\mathcal{P}$, then the $\mathcal{P}$-targeting problem has a solution. For example, if $\mathcal{P}$ is the property that $A$ is invertible, then $B$ must be invertible and hence the data must ensure that $B_{1}$ has full rank. The following corollary reveals a familiar criterion for row equivalence of two matrices; see [3, Theorem 4.4.1 and P.8.30].
Corollary 2.3.
Let $X,Y\in\mathsf{M}_{m\times n}(\mathbb{F})$ with $m\geq n\geq 1$ and $X\neq 0$. There is an invertible $A\in\mathsf{M}_{m}(\mathbb{F})$ such that $AX=Y$ if and only if $\operatorname{null}X=\operatorname{null}Y$.
Proof.
($\Rightarrow$) If $Y=AX$ and $A$ is invertible, then $X=A^{-1}Y$. Consequently, $\operatorname{null}X\subseteq\operatorname{null}Y$ and $\operatorname{null}Y\subseteq\operatorname{null}X$, so $\operatorname{null}X=\operatorname{null}Y$.
($\Leftarrow$) Let $r=\operatorname{rank}X\geq 1$. There is a choice of $B_{2}$ in (4) that makes $B$ invertible if and only if $\operatorname{rank}B_{1}=r$. If $\operatorname{null}X=\operatorname{null}Y$, then $\operatorname{rank}X=\operatorname{rank}Y$, and the relations
[…]
ensure that $\operatorname{rank}B_{1}=r$. ∎
3. Two Lemmas
The following lemma memorializes some consequences of the basic assumption $\operatorname{null}X\subseteq\operatorname{null}Y$.
Lemma 3.1.
Let $X,Y\in\mathsf{M}_{m\times n}(\mathbb{F})$ with $m\geq n\geq 1$ and $X\neq 0$. Suppose that $\operatorname{null}X\subseteq\operatorname{null}Y$ and adopt the notation used in the proof of Theorem 2.2. Then $X^{*}Y$ is unitarily similar to $\Sigma_{r}Z_{1}\oplus 0_{n-r}$, in which $Z_{1}=V_{1}^{*}YW_{1}$.
Proof.
Compute
[…]
In our computations with block matrices, we require the three types of Schur complements in the next lemma.
Lemma 3.2.
Let $H\in\mathsf{M}_{r}(\mathbb{F})$, $L\in\mathsf{M}_{(m-r)\times r}(\mathbb{F})$, and $\lambda\in\mathbb{F}$. Let
$B=\begin{bmatrix}H&L^{*}\\ L&\lambda I_{m-r}\end{bmatrix}\in\mathsf{M}_{m}(\mathbb{F}).$
  1. (a)
    If $\0\lambda\in\mathbb{R}\backslash\{0\}$, then $B$ is congruent to $(H-\lambda^{-1}L^{*}L)\oplus(\lambda I_{m-r})$.
  2. (b)
    If $H$ is Hermitian and invertible, then $B$ is congruent to $H\oplus(\lambda I_{m-r}-LH^{-1}L^{*})$.
  3. (c)
    If $H$ is Hermitian and $\operatorname{null}H\subseteq\operatorname{null}L$, then $B$ is congruent to $H\oplus(\lambda I_{m-r}-LH^{\dagger}L^{*})$.
Proof.
(a) Compute the congruence
$\begin{bmatrix}I_{r}&-\lambda^{-1}L^{*}\\ 0&I_{m-r}\end{bmatrix}\begin{bmatrix}H&L^{*}\\ L&\lambda I_{m-r}\end{bmatrix}\begin{bmatrix}I_{r}&0\\ -\lambda^{-1}L&I_{m-r}\end{bmatrix}=\begin{bmatrix}H-\lambda^{-1}L^{*}L&0\\ 0&\lambda I_{m-r}\end{bmatrix}.$
(b) The product
$\begin{bmatrix}I_{r}&0\\ -LH^{-1}&I_{m-r}\end{bmatrix}\begin{bmatrix}H&L^{*}\\ L&\lambda I_{m-r}\end{bmatrix}\begin{bmatrix}I_{r}&-H^{-1}L^{*}\\ 0&I_{m-r}\end{bmatrix}=\begin{bmatrix}H&0\\ 0&\lambda I_{m-r}-LH^{-1}L^{*}\end{bmatrix}$
is a congruence since $H$ is Hermitian.
(c) Since $I-H^{\dagger}H$ is the orthogonal projection onto $\operatorname{null}H$ and $\operatorname{null}H\subseteq\operatorname{null}L$, we have $L(I_{m-r}-H^{\dagger}H)=0$ and the product
$\begin{bmatrix}I_{r}&0\\ -LH^{\dagger}&I_{m-r}\end{bmatrix}\begin{bmatrix}H&L^{*}\\ L&\lambda I_{m-r}\end{bmatrix}\begin{bmatrix}I_{r}&-H^{\dagger}L^{*}\\ 0&I_{m-r}\end{bmatrix}=\begin{bmatrix}H&0\\ 0&\lambda I_{m-r}-LH^{\dagger}L^{*}\end{bmatrix}$
is a congruence. ∎
4. Hermitian targeting
Theorem 4.1.
Let $X,Y\in\mathsf{M}_{m\times n}(\mathbb{F})$ with $m\geq n\geq 1$ and $X\neq 0$. There is a Hermitian $A\in\mathsf{M}_{m}(\mathbb{F})$ such that $AX=Y$ if and only if $\operatorname{null}X\subseteq\operatorname{null}Y$ and $X^{*}Y$ is Hermitian.
Proof.
($\Rightarrow$) If $AX=Y$ and $A$ is Hermitian, then $\operatorname{null}X\subseteq\operatorname{null}Y$ and
$(X^{*}Y)^{*}=(X^{*}AX)^{*}=X^{*}AX=X^{*}Y.$
($\Leftarrow$) Let $r=\operatorname{rank}X$. It suffices to show that there is a choice of $B_{2}$ in (4) such that $B$ is Hermitian. Let
$B=\begin{bmatrix}Z_{1}\Sigma_{r}^{-1}&(Z_{2}\Sigma_{r}^{-1})^{*}\\ Z_{2}\Sigma_{r}^{-1}&\lambda I_{m-r}\end{bmatrix}=\begin{bmatrix}B_{1}&B_{2}\end{bmatrix},$ (7)
in which $\lambda\in\mathbb{R}$. Since $X^{*}Y$ is Hermitian, (6) ensures that $\Sigma_{r}Z_{1}$ is Hermitian, which implies that $\Sigma_{r}^{-*}(\Sigma_{r}Z_{1})\Sigma_{r}^{-1}=Z_{1}\Sigma_{r}^{-1}$ is Hermitian and hence $B$ is Hermitian. ∎
Corollary 4.2.
Let $X,Y\in\mathsf{M}_{m\times n}(\mathbb{F})$ with $m\geq n\geq 1$ and $X\neq 0$. There is an invertible Hermitian $A\in\mathsf{M}_{m}(\mathbb{F})$ such that $AX=Y$ if and only if $\operatorname{null}X=\operatorname{null}Y$ and $X^{*}Y$ is Hermitian.
Proof.
($\Rightarrow$) It follows from Theorem 4.1 that $X^{*}Y$ is Hermitian and $\operatorname{null}X\subseteq\operatorname{null}Y$. Since $X=A^{-1}Y$, it follows that $\operatorname{null}Y\subseteq\operatorname{null}X$, so $\operatorname{null}X=\operatorname{null}Y$.
($\Leftarrow$) Let $r=\operatorname{rank}X$. The rank-nullity theorem ensures that $\operatorname{rank}Y=r$ since $\dim\operatorname{null}X=\dim\operatorname{null}Y$. Let
$B=[B_{1}\,\,B_{2}]=\begin{bmatrix}Z_{1}\Sigma_{r}^{-1}&(Z_{2}\Sigma_{r}^{-1})^{*}\\ Z_{2}\Sigma_{r}^{-1}&\lambda I_{m-r}\end{bmatrix}=\begin{bmatrix}H&L^{*}\\ L&\lambda I_{m-r}\end{bmatrix}$
be the Hermitian matrix defined in (7), in which $H=Z_{1}\Sigma_{r}^{-1}\in\mathsf{M}_{r}(\mathbb{F})$ is Hermitian, $L=Z_{2}\Sigma_{r}^{-1}\in\mathsf{M}_{(m-r)\times r}(\mathbb{F})$, $B_{1}\in\mathsf{M}_{m\times r}(\mathbb{F})$, and $\lambda\in\mathbb{R}$ is nonzero. Then (5) ensures that $B_{1}$ has full column rank, and hence
$\operatorname{null}H\cap\operatorname{null}(L^{*}L)=\operatorname{null}H\cap\operatorname{null}L=\{\mathbf{0}\}.$ (8)
Lemma 3.2.a implies that
$\det B=(\det(\lambda I_{m-r}))(\det(H-\lambda^{-1}L^{*}L))=\lambda^{m-r}\det(H-\lambda^{-1}L^{*}L).$
We need to show that there is some real nonzero $\lambda$ such that $\det(H-\lambda^{-1}L^{*}L)\neq 0$. Consider the analytic function $f(z)=\det(H-z^{-1}L^{*}L)$ on \$\0\mathbb{C}\backslash\{0\}$. If $f(\lambda)=0$ for every real $\lambda\neq 0$, then $f(z)=0$ for every $\0z\in\mathbb{C}\backslash\{0\}$. To show that $f(\lambda)\neq 0$ for some real nonzero $\lambda$, it suffices to show that $f(i)\neq 0$. If $f(i)=0$, there is a nonzero $\mathbf{u}\in\mathbb{C}^{r}$ such that $(H+iL^{*}L)\mathbf{u}=\mathbf{0}$. However,
$(H+iL^{*}L)\mathbf{u}=\mathbf{0}$ $\implies\mathbf{u}^{*}H\mathbf{u}+i\mathbf{u}^{*}L^{*}L\mathbf{u}=\mathbf{u}^{*}(H+iL^{*}L)\mathbf{u}=0$
$\implies\operatorname{Im}(\mathbf{u}^{*}H\mathbf{u}+i\mathbf{u}^{*}L^{*}L\mathbf{u})=0$
$\implies(L\mathbf{u})^{*}(L\mathbf{u})=\mathbf{u}^{*}L^{*}L\mathbf{u}=0$
$\implies L\mathbf{u}=\mathbf{0}$
$\implies H\mathbf{u}=(H+iL^{*}L)\mathbf{u}=\mathbf{0}$
$\implies\operatorname{null}H\cap\operatorname{null}L\neq\{\mathbf{0}\},$
which contradicts (8). We conclude that $B$ is invertible and Hermitian for some real nonzero $\lambda$. ∎
If a nonzero $Y\in\mathsf{M}_{m\times n}(\mathbb{F})$ is given, what are all the $X\in\mathsf{M}_{m\times n}(\mathbb{F})$ such that $Y=AX$ for some Hermitian $A?$ To analyze this question we start with a singular value decomposition of $Y$ instead of $X$.
Corollary 4.3.
Let $X,Y\in\mathsf{M}_{m\times n}(\mathbb{F})$ with $m\geq n\geq 1$ and $r=\operatorname{rank}Y\geq 1$. Let $Y=V\Sigma W^{*}$ be a singular value decomposition, in which
$V^{*}YW=\Sigma=\begin{bmatrix}\Sigma_{r}&0\\ 0&0\end{bmatrix}\in\mathsf{M}_{m\times n}(\mathbb{R})$
and $\Sigma_{r}\in\mathsf{M}_{r}(\mathbb{R})$ is diagonal and positive definite. Let $Z=V^{*}XW$ and partition
$Z=\begin{bmatrix}Z_{11}&Z_{12}\\ Z_{21}&Z_{22}\end{bmatrix}$
conformally to $\Sigma$. Let $Z_{1}=\Big{[}\begin{smallmatrix}Z_{11}\\ Z_{21}\end{smallmatrix}\Big{]}\in\mathsf{M}_{m\times r}(\mathbb{F})$.
  1. (a)
    $X^{*}Y=Y^{*}X$ if and only if $Z_{12}=0$ and $\Sigma_{r}Z_{11}=Z_{11}^{*}\Sigma_{r}$.
  2. (b)
    If $Z_{12}=0$, $\Sigma_{r}Z_{11}=Z_{11}^{*}\Sigma_{r}$, $\operatorname{rank}Z_{1}=r$, and $(Z_{21}\operatorname{null}Z_{11})\cap\operatorname{col}Z_{22}=\{\mathbf{0}\}$, then there is a Hermitian $A\in\mathsf{M}_{m}(\mathbb{F})$ such that $AX=Y$.
  3. (c)
    If $Z_{12}=0$, $\Sigma_{r}Z_{11}=Z_{11}^{*}\Sigma_{r}$, and $Z_{11}$ is invertible, then for each choice of $Z_{21}\in\mathsf{M}_{(m-r)\times r}(\mathbb{F})$ and $Z_{22}\in\mathsf{M}_{(m-r)\times(n-r)}(\mathbb{F})$ there is a Hermitian $A\in\mathsf{M}_{m}(\mathbb{F})$ such that $AX=Y$.
Proof.
(a) A computation reveals that $X^{*}Y$ is Hermitian if and only if
$X^{*}Y=Y^{*}X$ $\iff X^{*}V\Sigma W^{*}=W\Sigma^{\mathsf{T}}V^{*}X$
$\iff(W^{*}X^{*}V)\Sigma=\Sigma^{\mathsf{T}}(V^{*}XW)$
$\iff\begin{bmatrix}Z_{11}^{*}&Z_{21}^{*}\\ Z_{12}^{*}&Z_{22}^{*}\end{bmatrix}\begin{bmatrix}\Sigma_{r}&0\\ 0&0\end{bmatrix}=\begin{bmatrix}\Sigma_{r}&0\\ 0&0\end{bmatrix}\begin{bmatrix}Z_{11}&Z_{12}\\ Z_{21}&Z_{22}\end{bmatrix}$
$\iff\begin{bmatrix}Z_{11}^{*}\Sigma_{r}&0\\ Z_{12}^{*}\Sigma_{r}&0\end{bmatrix}=\begin{bmatrix}\Sigma_{r}Z_{11}&\Sigma_{r}Z_{12}\\ 0&0\end{bmatrix}.$
Therefore, $X^{*}Y$ is Hermitian if and only if $Z_{12}=0$ and $\Sigma_{r}Z_{11}=Z_{11}^{*}\Sigma_{r}$.
(b) We claim that $\operatorname{null}X\subseteq\operatorname{null}Y$ if and only if $\operatorname{null}Z\subseteq\operatorname{null}\Sigma$. To validate this claim, suppose that $\operatorname{null}X\subseteq\operatorname{null}Y$. Then
$\mathbf{u}\in\operatorname{null}Z$ $\implies\mathbf{0}=Z\mathbf{u}=(V^{*}XW)\mathbf{u}=\mathbf{0}$
$\implies X(W\mathbf{u})=\mathbf{0}$
$\implies Y(W\mathbf{u})=\mathbf{0}$
$\implies V^{*}Y(W\mathbf{u})=\mathbf{0}$
$\implies\mathbf{0}=(V^{*}YW)\mathbf{u}=\Sigma\mathbf{u}$
$\implies\mathbf{u}\in\operatorname{null}\Sigma,$
from which we conclude that $\operatorname{null}Z\subseteq\operatorname{null}\Sigma$. The converse follows from a parallel argument. Now assume that $Z_{12}=0$ and $\Sigma_{r}Z_{11}=Z_{11}^{*}\Sigma_{r}$, which ensures that $X^{*}Y=Y^{*}X$. Let $\mathbf{u}=\big{[}\begin{smallmatrix}\mathbf{u}_{1}\\ \mathbf{u}_{2}\end{smallmatrix}\big{]}\in\mathbb{F}^{n}$, in which $\mathbf{u}_{1}\in\mathbb{F}^{r}$. Then $\mathbf{u}\in\operatorname{null}\Sigma$ if and only if $\mathbf{u}_{1}=\mathbf{0}$. If $\operatorname{rank}Z_{1}=r$ and there is a $\mathbf{u}\in\operatorname{null}Z$ such that $\mathbf{u}_{1}\neq\mathbf{0}$, then
$\mathbf{0}=Z\mathbf{u}=\begin{bmatrix}Z_{11}&0\\ Z_{21}&Z_{22}\end{bmatrix}\begin{bmatrix}\mathbf{u}_{1}\\ \mathbf{u}_{2}\end{bmatrix}=\begin{bmatrix}Z_{11}\mathbf{u}_{1}\\ Z_{21}\mathbf{u}_{1}+Z_{22}\mathbf{u}_{2}\end{bmatrix}=\begin{bmatrix}\mathbf{0}\\ Z_{21}\mathbf{u}_{1}+Z_{22}\mathbf{u}_{2}\end{bmatrix}.$
Since $\operatorname{rank}Z_{1}=r$, it follows that $Z_{21}\mathbf{u}_{1}\neq\mathbf{0}$, which implies that $Z_{22}\mathbf{u}_{2}\neq\mathbf{0}$ and hence $(Z_{21}\operatorname{null}Z_{11})\cap\operatorname{col}Z_{22}\neq\{\mathbf{0}\}$. We conclude that $\operatorname{null}Z\subseteq\operatorname{null}\Sigma$.
(c) If $Z_{11}$ is invertible, then $\operatorname{rank}Z_{1}=r$ and $\operatorname{null}Z_{11}=\{\mathbf{0}\}$, so $(Z_{21}\operatorname{null}Z_{11})\cap\operatorname{col}Z_{22}\subseteq Z_{21}\operatorname{null}Z_{11}=\{\mathbf{0}\}$. ∎
5. Positive semidefinite targeting
We now investigate the positive semidefinite and positive definite targeting problems. An important property of positive semidefinite matrices $A$ is that $A\mathbf{u}=\mathbf{0}$ if and only if $\mathbf{u}^{*}A\mathbf{u}=0$; see [3, Corollary 15.1.18] or [4, Observation 7.1.6].
Theorem 5.1.
Let $X,Y\in\mathsf{M}_{m\times n}(\mathbb{F})$ with $m\geq n\geq 1$ and $X\neq 0$. There is a positive semidefinite $A\in\mathsf{M}_{m}(\mathbb{F})$ such that $AX=Y$ if and only if $X^{*}Y\geq 0$ and $\operatorname{null}Y=\operatorname{null}(X^{*}Y)$.
Proof.
($\Rightarrow$) If $AX=Y$ and $A\geq 0$, then $X^{*}Y=X^{*}AX$, which is positive semidefinite. To verify the assertion about null spaces, use the fact that $X^{*}Y\geq 0$ and observe that
$\mathbf{u}\in\operatorname{null}(X^{*}Y)$ $\implies(X^{*}Y)\mathbf{u}=\mathbf{0}$
$\implies\mathbf{u}^{*}(X^{*}Y)\mathbf{u}=0$
$\implies\mathbf{u}^{*}(X^{*}AX)\mathbf{u}=0$
$\implies(X\mathbf{u})^{*}A(X\mathbf{u})=0$
$\implies A(X\mathbf{u})=\mathbf{0}$
$\implies(AX)\mathbf{u}=\mathbf{0}$
$\implies Y\mathbf{u}=\mathbf{0}$
$\implies\mathbf{u}\in\operatorname{null}Y.$
($\Leftarrow$) Let $r=\operatorname{rank}X$. Since $X^{*}Y$ is Hermitian, we have $\operatorname{null}X\subseteq\operatorname{null}(Y^{*}X)=\operatorname{null}(X^{*}Y)=\operatorname{null}Y$, so $\operatorname{null}X\subseteq\operatorname{null}Y$. The hypotheses of Theorem 4.1 are satisfied, and its proof shows that
$B=\begin{bmatrix}B_{1}&B_{2}\end{bmatrix}=\begin{bmatrix}Z_{1}\Sigma_{r}^{-1}&(Z_{2}\Sigma_{r}^{-1})^{*}\\ Z_{2}\Sigma_{r}^{-1}&\lambda I_{m-r}\end{bmatrix}=\begin{bmatrix}H&L^{*}\\ L&\lambda I_{m-r}\end{bmatrix}$ (9)
is Hermitian for any $\lambda>0$. The identity (6) and the assumption $X^{*}Y\geq 0$ imply that $\Sigma_{r}Z_{1}\geq 0$, so $\Sigma_{r}^{-*}(\Sigma_{r}Z_{1})\Sigma_{r}^{-1}=Z_{1}\Sigma_{r}^{-1}=H\geq 0$. The hypothesis $\operatorname{null}Y=\operatorname{null}(X^{*}Y)$, the fact that $H\geq 0$, and (5) ensure that
$\operatorname{null}H=\operatorname{null}B_{1}=\operatorname{null}H\cap\operatorname{null}L\subseteq\operatorname{null}L.$ (10)
Let $\lambda_{1}\geq 0$ be the largest eigenvalue of the positive semidefinite matrix $LH^{\dagger}L^{*}$. We claim that $B\geq 0$ if $\lambda\geq\lambda_{1}$. Lemma 3.2.c and (10) ensure that $B$ is congruent to $H\oplus(\lambda I_{m-r}-LH^{\dagger}L^{*})$, which is positive semidefinite if $\lambda I_{m-r}-LH^{\dagger}L^{*}\geq 0$. If $\lambda\geq\lambda_{1}$ and $\mathbf{x}\in\mathbb{F}^{m-r}$, then
$\mathbf{x}^{*}(\lambda I_{m-r}-LH^{\dagger}L^{*})\mathbf{x}=\lambda\mathbf{x}^{*}\mathbf{x}-\mathbf{x}^{*}(LH^{\dagger}L^{*})\mathbf{x}\geq(\lambda-\lambda_{1})\|\mathbf{x}\|^{2}\geq 0,$
which validates our claim. ∎
Corollary 5.2.
Let $X,Y\in\mathsf{M}_{m\times n}(\mathbb{F})$ with $m\geq n\geq 1$ and $X\neq 0$. There is a positive definite $A\in\mathsf{M}_{m}(\mathbb{F})$ such that $AX=Y$ if and only if $X^{*}Y\geq 0$, $\operatorname{null}Y=\operatorname{null}(X^{*}Y)$, and $\operatorname{rank}Y=\operatorname{rank}X$.
Proof.
($\Rightarrow$) If $A\in\mathsf{M}_{m}(\mathbb{F})$ is positive definite and $AX=Y$, then $A$ is invertible and $\operatorname{rank}X=\operatorname{rank}Y$. Moreover, $\operatorname{null}Y=\operatorname{null}(X^{*}Y)$ and $X^{*}Y\geq 0$, as in the preceding theorem.
($\Leftarrow$) Let $r=\operatorname{rank}X$. Since $\operatorname{rank}Y=\operatorname{rank}X$ and $\operatorname{null}Y=\operatorname{null}(X^{*}Y)$, it follows that
$r=\operatorname{rank}Y=\operatorname{rank}(X^{*}Y)=\operatorname{rank}Z_{1}=\operatorname{rank}H.$
Thus, $H\geq 0$ (as in the preceding theorem), $H\in\mathsf{M}_{r}(\mathbb{F})$ is invertible, and $H$ is positive definite. Lemma 3.2.b ensures that $B$ in (9) is congruent to $H\oplus(\lambda I_{n-r}-LH^{-1}L^{*})$, which is positive definite if $\lambda$ is greater than the largest eigenvalue of $LH^{-1}L^{*}$. ∎
Corollary 5.3.
Let $X,Y\in\mathsf{M}_{m\times n}(\mathbb{F})$ with $m\geq n\geq 1$ and $\operatorname{rank}X=n$. There is a positive definite $A\in\mathsf{M}_{m}$ such that $AX=Y$ if and only if $X^{*}Y>0$.
Proof.
($\Rightarrow$) If $A\in\mathsf{M}_{m}(\mathbb{F})$ is positive definite and $AX=Y$, then $X^{*}Y=X^{*}AX$, which is positive definite since $X$ has full rank.
($\Leftarrow$) If $X^{*}Y>0$, then $n=\operatorname{rank}(X^{*}Y)\leq\operatorname{rank}Y$, so $\operatorname{rank}Y=n$ and $\operatorname{null}Y=\{\mathbf{0}\}=\operatorname{null}(X^{*}Y)$. The preceding corollary ensures that there is a positive definite $A\in\mathsf{M}_{m}$ such that $AX=Y$. ∎
6. Unitary targeting
The unitary targeting problem generalizes the familiar fact that, for given $\mathbf{x},\mathbf{y}\in\mathbb{F}^{n}$, there is a unitary $U\in\mathsf{M}_{n}(\mathbb{F})$ such that $\mathbf{y}=U\mathbf{x}$ if and only if $\|\mathbf{x}\|=\|\mathbf{y}\|$.
Theorem 6.1.
Let $X,Y\in\mathsf{M}_{m\times n}(\mathbb{F})$ with $m\geq n\geq 1$ and $X\neq 0$. There is a unitary $A\in\mathsf{M}_{m}(\mathbb{F})$ such that $AX=Y$ if and only if $X^{*}X=Y^{*}Y$.
Proof.
($\Rightarrow$) If $Y=AX$ and $A$ is unitary, then $Y^{*}Y=(AX)^{*}(AX)=X^{*}(A^{*}A)X=X^{*}IX=X^{*}X$.
($\Leftarrow$) Let $r=\operatorname{rank}X$. If $X^{*}X=Y^{*}Y$, then $\operatorname{null}X=\operatorname{null}Y$ [3, Theorem 15.1.9] and (3) ensures that $B=\begin{bmatrix}B_{1}&B_{2}\end{bmatrix}$ satisfies (2) with $B_{1}=Z\Sigma_{r}^{-1}$. It suffices to show that $B_{1}$ has orthonormal columns, since we may then choose $B_{2}$ to make $B$ unitary; see [3, Corollary 6.3.14]. Observe that
[…]
We also have
$W^{*}(X^{*}X)W=\Sigma^{*}\Sigma=\begin{bmatrix}\Sigma_{r}^{2}&0\\ 0&0\end{bmatrix}.$
Since $X^{*}X=Y^{*}Y$, we conclude that $Z^{*}Z=\Sigma_{r}^{2}$ and hence
$I_{r}=(Z\Sigma_{r}^{-1})^{*}(Z\Sigma_{r}^{-1})=B_{1}^{*}B_{1},$
that is, $B_{1}$ has orthonormal columns. ∎
The polar decomposition [3, § 16.3] or [4, § 7.3] provides an alternative approach to the unitary targeting problem. If $X^{*}X=Y^{*}Y$ and $Q=(X^{*}X)^{1/2}$, then $X=U_{1}Q$ and $Y=V_{1}Q$, in which $U_{1},V_{1}\in\mathsf{M}_{m\times n}(\mathbb{F})$ have orthonormal columns. If $U=[U_{1}\,\,U_{2}]$ and $V=[V_{1}\,\,V_{2}]$ are unitary, then $VU^{*}$ is unitary and
$(VU^{*})X=(VU^{*})U_{1}Q=V(U^{*}U_{1})Q=[V_{1}\,\,V_{2}]\begin{bmatrix}I\\ 0\end{bmatrix}Q=V_{1}Q=Y.$
If $X$ and $Y$ do not have the same numbers of columns, there is an interesting story to be told about the identity $X^{*}X=Y^{*}Y$. For details, examples, and a historical review, see [5].
7. Reflection Targeting
A matrix $A\in\mathsf{M}_{m}(\mathbb{F})$ is a reflection if it is a Hermitian involution ($A=A^{*}$ and $A^{2}=I)$, or, equivalently, if it is Hermitian and unitary. We now consider the reflection targeting problem.
Theorem 7.1.
Let $X,Y\in\mathsf{M}_{m\times n}(\mathbb{F})$ with $m\geq n\geq 1$ and $r=\operatorname{rank}Y\geq 1$. There is a reflection $A\in\mathsf{M}_{m}(\mathbb{F})$ such that $AX=Y$ if and only if $X^{*}Y$ is Hermitian and $X^{*}X=Y^{*}Y$.
Proof.
($\Rightarrow$) If $A$ is a reflection and $AX=Y$, then $A$ is Hermitian and so is $X^{*}Y=X^{*}AX$. Since $A$ is unitary, $Y^{*}Y=(AX)^{*}AX=X^{*}A^{*}AX=X^{*}IX=X^{*}X$.
($\Leftarrow$) Suppose that $X^{*}Y=Y^{*}X$ and $X^{*}X=Y^{*}Y$. Let $E=X+Y$ and $F=X-Y$, so that $X=\frac{1}{2}(E+F)$ and $Y=\frac{1}{2}(E-F)$. Let $P=F(F^{*}F)^{\dagger}F^{*}$, which is the (necessarily Hermitian) orthogonal projection onto $\operatorname{col}F$. Define the Hermitian matrix $A=I-2P$, which is a reflection since
$A^{2}=(I-2P)^{2}=I-4P+4P^{2}=I-4P+4P=I.$
Observe that $\operatorname{col}E$ is orthogonal to $\operatorname{col}F$ since
$E^{*}F=(X+Y)^{*}(X-Y)=(X^{*}X-Y^{*}Y)+(Y^{*}X-X^{*}Y)=0+0=0.$
Therefore, $PE=0$ and $PF=F$. Consequently,
$AX$ $=\frac{1}{2}(I-2P)(E+F)=\frac{1}{2}(E+F-2PE-2PF)$
$=\frac{1}{2}(E+F-0-2F)=\frac{1}{2}(E-F)=Y.\qed$
For a given nonzero $Y\in\mathsf{M}_{m\times n}(\mathbb{F})$, the following corollary shows how to construct all the $X\in\mathsf{M}_{m\times n}(\mathbb{F})$ such that $Y=AX$ for some reflection $A$.
Corollary 7.2.
Let $X,Y\in\mathsf{M}_{m\times n}(\mathbb{F})$ with $m\geq n\geq 1$ and $r=\operatorname{rank}Y\geq 1$. Let $Y=V\Sigma W^{*}$ be a singular value decomposition, in which
$V^{*}YW=\Sigma=\begin{bmatrix}\Sigma_{r}&0\\ 0&0\end{bmatrix}\in\mathsf{M}_{m\times n}(\mathbb{R})$
and $\Sigma_{r}\in\mathsf{M}_{r}(\mathbb{R})$ is diagonal and positive definite. There is a reflection $A\in\mathsf{M}_{m}(\mathbb{F})$ such that $AX=Y$ if and only if
$V^{*}XW=\begin{bmatrix}U_{11}\Sigma_{r}&0\\ U_{21}\Sigma_{r}&0\end{bmatrix},$ (11)
in which $U_{11}\in\mathsf{M}_{r}(\mathbb{F})$ is Hermitian and $U_{1}=\big{[}\begin{smallmatrix}U_{11}\\ U_{21}\end{smallmatrix}\big{]}\in\mathsf{M}_{m\times r}(\mathbb{F})$ has orthonormal columns.
Proof.
($\Rightarrow$) Partition
$V^{*}XW=\begin{bmatrix}Z_{11}&Z_{12}\\ Z_{21}&Z_{22}\end{bmatrix}$
conformally to $\Sigma$. Since $X^{*}Y$ is Hermitian, Corollary 4.3 ensures that $Z_{12}=0$ and $\Sigma_{r}Z_{11}=Z_{11}^{*}\Sigma_{r}$. Since $X^{*}X=Y^{*}Y$, we have
$W^{*}X^{*}XW$ $=\begin{bmatrix}Z_{11}^{*}&Z_{21}^{*}\\ 0&Z_{22}^{*}\end{bmatrix}\begin{bmatrix}Z_{11}&0\\ Z_{21}&Z_{22}\end{bmatrix}$
$=\begin{bmatrix}Z_{11}^{*}Z_{11}+Z_{21}^{*}Z_{21}&Z_{21}^{*}Z_{22}\\ Z_{22}^{*}Z_{21}&Z_{22}^{*}Z_{22}\end{bmatrix}=\begin{bmatrix}\Sigma_{r}^{2}&0\\ 0&0\end{bmatrix}$ (12)
$=W^{*}Y^{*}YW.$
A comparison of entries in (12) reveals that $Z_{22}^{*}Z_{22}=0$, and hence $Z_{22}=0$. If we partition
[…]
another comparison of entries shows that $Z_{1}^{*}Z_{1}=\Sigma_{r}^{2}$. It follows from the polar decomposition that $Z_{1}=U_{1}\Sigma_{r}$, in which $U_{1}=\big{[}\begin{smallmatrix}U_{11}\\ U_{21}\end{smallmatrix}\big{]}\in\mathsf{M}_{m\times r}(\mathbb{F})$ has orthonormal columns. The condition $\Sigma_{r}Z_{11}=Z_{11}^{*}\Sigma_{r}$ ensures that $\Sigma_{r}U_{11}\Sigma_{r}=\Sigma_{r}U_{11}^{*}\Sigma_{r}$, so $U_{11}$ is Hermitian.
($\Leftarrow$) If $V^{*}XW$ can be partitioned as in (11), then
$W^{*}(X^{*}X)W$ $=\begin{bmatrix}\Sigma_{r}U_{1}^{*}\\ 0\end{bmatrix}\begin{bmatrix}U_{1}\Sigma_{r}&0\end{bmatrix}=\begin{bmatrix}\Sigma_{r}U_{1}^{*}U_{1}\Sigma_{r}&0\\ 0&0\end{bmatrix}$
$=\begin{bmatrix}\Sigma_{r}^{2}&0\\ 0&0\end{bmatrix}=W^{*}(Y^{*}Y)W,$
which shows that $X^{*}X=Y^{*}Y$. Also,
$W^{*}(X^{*}Y)W$ $=\begin{bmatrix}\Sigma_{r}U_{11}^{*}&\Sigma_{r}U_{21}^{*}\\ 0&0\end{bmatrix}\begin{bmatrix}\Sigma_{r}&0\\ 0&0\end{bmatrix}$
$=\begin{bmatrix}\Sigma_{r}U_{11}^{*}\Sigma_{r}&0\\ 0&0\end{bmatrix}=\begin{bmatrix}\Sigma_{r}U_{11}\Sigma_{r}&0\\ 0&0\end{bmatrix}$
is Hermitian, which shows that $X^{*}Y$ is Hermitian. The preceding theorem now ensures that there is a reflection $A\in\mathsf{M}_{m}(\mathbb{F})$ such that $AX=Y$. ∎
One consequence of Theorem 7.1 is that, for given $\mathbf{x},\mathbf{y}\in\mathbb{C}^{n}$, there is a reflection $A\in\mathsf{M}_{n}(\mathbb{C})$ such that $A\mathbf{x}=\mathbf{y}$ if and only if $\|\mathbf{x}\|=\|\mathbf{y}\|$ and $\mathbf{x}^{*}\mathbf{y}$ is real; the latter requirement is superfluous if $\mathbb{F}=\mathbb{R}$. In fact, $A$ may be chosen to be a scalar multiple of a Householder matrix in this case. Our proof of Theorem 7.1 constructs a reflection matrix that is a natural analog of a Householder matrix.
8. Orthogonal Projection Targeting
A computation with the Moore–Penrose pseudoinverse provides a necessary and sufficient condition for there to be a solution to the orthogonal projection targeting problem.
Theorem 8.1.
Let $X,Y\in\mathsf{M}_{m\times n}(\mathbb{F})$ with $m\geq n\geq 1$ and $Y\neq 0$. There is an orthogonal projection $A\in\mathsf{M}_{m}(\mathbb{F})$ such that $AX=Y$ if and only if $Y^{*}X=Y^{*}Y$.
Proof.
($\Rightarrow$) If $Y=AX$, $A$ is Hermitian, and $A^{2}=A$, then
$Y^{*}Y=(AX)^{*}(AX)=X^{*}A^{*}AX=X^{*}A^{2}X=X^{*}(AX)=X^{*}Y.$
($\Leftarrow$) If $Y^{*}X=Y^{*}Y$, let $A=Y(Y^{*}Y)^{\dagger}Y^{*}$. Then $A$ is Hermitian and
$A^{2}=Y(Y^{*}Y)^{\dagger}(Y^{*}Y)(Y^{*}Y)^{\dagger}Y^{*}=Y(Y^{*}Y)^{\dagger}Y^{*}=A,$
so $A$ is an orthogonal projection. Since $(Y^{*}Y)^{\dagger}(Y^{*}Y)$ is the orthogonal projection on $\operatorname{col}(Y^{*}Y)^{*}=\operatorname{col}(Y^{*}Y)=\operatorname{col}Y^{*}$, it follows that $(Y^{*}Y)^{\dagger}(Y^{*}Y)Y^{*}=Y^{*}$ and hence $Y=Y(Y^{*}Y)(Y^{*}Y)^{\dagger}$. Therefore,
$AX=Y(Y^{*}Y)^{\dagger}(Y^{*}X)=Y(Y^{*}Y)^{\dagger}(Y^{*}Y)=Y.\qed$
We can use a singular value decomposition of $Y$ (note the change of strategy here) to identify all solutions $X$ to the orthogonal projection targeting problem $AX=Y$ for a given $Y$.
Corollary 8.2.
Let $X,Y\in\mathsf{M}_{m\times n}(\mathbb{F})$ with $m\geq n\geq 1$ and $r=\operatorname{rank}Y\geq 1$. Let $Y=V\Sigma W^{*}$ be a singular value decomposition, in which
$V^{*}YW=\Sigma=\begin{bmatrix}\Sigma_{r}&0\\ 0&0\end{bmatrix}\in\mathsf{M}_{m\times n}(\mathbb{R})$
and $\Sigma_{r}\in\mathsf{M}_{r}(\mathbb{R})$ is diagonal and positive definite. Partition
$V^{*}XW=Z=\begin{bmatrix}Z_{11}&Z_{12}\\ Z_{21}&Z_{22}\end{bmatrix}$
conformally to $\Sigma$. There is an orthogonal projection $A\in\mathsf{M}_{m}(\mathbb{F})$ such that $AX=Y$ if and only if
$V^{*}XW=\begin{bmatrix}\Sigma_{r}&0\\ Z_{21}&Z_{22}\end{bmatrix},$
in which $Z_{21}\in\mathsf{M}_{(m-r)\times r}(\mathbb{F})$ and $Z_{22}\in\mathsf{M}_{(m-r)\times(n-r)}(\mathbb{F})$.
Proof.
Compare entries in
$Y^{*}Y=Y^{*}X$ $\iff W\Sigma^{\mathsf{T}}\Sigma W^{*}=W\Sigma^{\mathsf{T}}V^{*}VZW^{*}$
$\iff\Sigma^{\mathsf{T}}\Sigma=\Sigma^{\mathsf{T}}Z$
$\iff\begin{bmatrix}\Sigma_{r}^{2}&0\\ 0&0\end{bmatrix}=\begin{bmatrix}\Sigma_{r}Z_{11}&\Sigma_{r}Z_{12}\\ 0&0\end{bmatrix}.\qed$
9. Complex symmetric targeting
A real symmetric matrix is Hermitian, but a complex symmetric matrix need not even be normal. However, there is an analog of Theorem 4.1 for complex symmetric targeting.
Theorem 9.1.
Let $X,Y\in\mathsf{M}_{m\times n}(\mathbb{C})$ with $m\geq n\geq 1$ and $X\neq 0$. There is a complex symmetric $A\in\mathsf{M}_{m}(\mathbb{C})$ such that $AX=Y$ if and only if $\operatorname{null}X\subseteq\operatorname{null}Y$ and $X^{\mathsf{T}}Y$ is complex symmetric.
Proof.
($\Rightarrow$) If $Y=AX$ and $A$ is complex symmetric, then $\operatorname{null}X\subseteq\operatorname{null}Y$, $X^{\mathsf{T}}AX$ is complex symmetric, and $X^{\mathsf{T}}Y=X^{\mathsf{T}}AX=(X^{\mathsf{T}}Y)^{\mathsf{T}}$.
($\Leftarrow$) Let $\operatorname{rank}X=r\geq 1$ and adopt the notation in the proof of Theorem 2.2. Let $X=V\Sigma W^{*}$ be a singular value decomposition. If there is a complex symmetric $F\in\mathsf{M}_{m}(\mathbb{C})$ such that $F\Sigma=V^{\mathsf{T}}YW$, then
$\overline{V}(F\Sigma)W^{*}=\overline{V}(V^{\mathsf{T}}YW)W^{*}=Y$
and
$Y=\overline{V}(F\Sigma)W^{*}=(\overline{V}FV^{*})(V\Sigma W^{*})=(\overline{V}FV^{*})X.$
Thus, $A=\overline{V}FV^{*}$ is complex symmetric and $AX=Y$. If $\operatorname{null}X\subseteq\operatorname{null}Y$, then $YW_{2}=0$ and
$V^{\mathsf{T}}YW=\begin{bmatrix}V^{\mathsf{T}}YW_{1}&0\end{bmatrix}.$
Partition $F=[F_{1}\,\,F_{2}]$ with
$F_{1}=\begin{bmatrix}V_{1}^{\mathsf{T}}YW_{1}\Sigma_{r}^{-1}\\[3.0pt] V_{2}^{\mathsf{T}}YW_{1}\Sigma_{r}^{-1}\end{bmatrix}\in\mathsf{M}_{m\times r}(\mathbb{C}).$
For any choice of $F_{2}\in\mathsf{M}_{m\times(m-r)}(\mathbb{C})$, we have $F\Sigma=V^{\mathsf{T}}YW$. Now compute
$W^{\mathsf{T}}(X^{\mathsf{T}}Y)W$ $=W^{\mathsf{T}}((V\Sigma W^{*})^{\mathsf{T}}Y)W=W^{\mathsf{T}}(\overline{W}\Sigma^{\mathsf{T}}V^{\mathsf{T}}Y)W$
$=\Sigma^{\mathsf{T}}(V^{\mathsf{T}}YW)=\Sigma^{\mathsf{T}}\begin{bmatrix}V^{\mathsf{T}}YW_{1}&0\end{bmatrix}$
$=\begin{bmatrix}\Sigma_{r}&0\\ 0&0\end{bmatrix}\begin{bmatrix}V_{1}^{\mathsf{T}}YW_{1}&0\\[3.0pt] V_{2}^{\mathsf{T}}YW_{1}&0\end{bmatrix}=\begin{bmatrix}\Sigma_{r}V_{1}^{\mathsf{T}}YW_{1}&0\\ 0&0\end{bmatrix}.$
If $X^{\mathsf{T}}Y$ is symmetric, then so are $\Sigma_{r}V_{1}^{\mathsf{T}}YW_{1}$ and $\Sigma_{r}^{-1}(\Sigma_{r}V_{1}^{\mathsf{T}}YW_{1})\Sigma_{r}^{-1}=V_{1}^{\mathsf{T}}YW_{1}\Sigma_{r}^{-1}$. If $G\in\mathsf{M}_{m-r}$ is any complex symmetric matrix, then
$F=\begin{bmatrix}V_{1}^{\mathsf{T}}YW_{1}\Sigma_{r}^{-1}&(V_{2}^{\mathsf{T}}YW_{1}\Sigma_{r}^{-1})^{\mathsf{T}}\\[3.0pt] V_{2}^{\mathsf{T}}YW_{1}\Sigma_{r}^{-1}&G\end{bmatrix}$
is complex symmetric and $F\Sigma=\begin{bmatrix}V^{\mathsf{T}}YW_{1}&0\end{bmatrix}=V^{\mathsf{T}}YW$. Thus, $A=\overline{V}FV^{*}$ is complex symmetric and $AX=Y$. ∎
10. Normal Targeting
[…]
For $n\geq 2$, however, it is not (yet) clear how to proceed. For a normal $A\in\mathsf{M}_{m}(\mathbb{F})$, we have $A^{*}\!A=AA^{*}$, which is a system of $m^{2}$ equations in the entries of $A$. The equations corresponding to diagonal entries of $A^{*}\!A$ and $AA^{*}$ say that, for each $i$, the $i$th row and the $i$th column of $A$ have the same Euclidean norm. If we partition $A=\big{[}\begin{smallmatrix}B&D\\ C&E\end{smallmatrix}\big{]}\in\mathsf{M}_{m}(\mathbb{F})$, in which $B\in\mathsf{M}_{k}(\mathbb{F})$, then inspection of the $(1,1)$ block of $A^{*}A=AA^{*}$ reveals that $B^{*}B-BB^{*}+C^{*}C=DD^{*}$. Consequently, for given $B$ and $C$, in order for there to be matrices $D$ and $E$ such that $A$ is normal, it is necessary that
$H(B,C)=B^{*}B-BB^{*}+C^{*}C$ (13)
be positive semidefinite. If $m=2k$, this necessary condition is sufficient for $k=2$ (see [1]), but not for $k\geq 3$. Consider
$B=\begin{bmatrix}0&0&1\\ 1&0&0\\ 0&1&0\end{bmatrix}\quad\text{and}\quad C=\begin{bmatrix}1&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix}.$
In this case, $B$ is normal and $H(B,C)=C$ is positive semidefinite. However, there is no choice of $D$ and $E$ that makes $A$ normal because the first row and first column of $A$ cannot have the same Euclidean norm. For a discussion of the condition (13) see [1] and [2].
Although the normal targeting problem remains open, the following theorem provides a solution in a special case that provides an alternative solution to the reflection and orthogonal projection targeting problems.
Theorem 10.1.
Let $X,Y\in\mathsf{M}_{m\times n}(\mathbb{C})$ and let $\lambda,\mu\in\mathbb{C}$ be distinct.
  1. (a)
    There is a normal $A\in\mathsf{M}_{m}(\mathbb{C})$ such that the spectrum of $A$ is $\{\lambda,\mu\}$ and $AX=Y$ if and only if $(Y-\lambda X)^{*}(Y-\mu X)=0$.
  2. (b)
    If $X$, $Y$, $\lambda$, and $\mu$ are real and $(Y-\lambda X)^{\mathsf{T}}(Y-\mu X)=0$, then there is a real normal $A\in\mathsf{M}_{m}(\mathbb{R})$ such that the spectrum of $A$ is $\{\lambda,\mu\}$ and $AX=Y$.
Proof.
(a) ()($\Rightarrow$)( ⇒ ) If $AX=Y$, $A$ is normal, and the eigenvalues of $A$ are $\lambda$ (with multiplicity $p\geq 1)$ and $\mu$ (with multiplicity $m-p\geq 1)$, then there is a unitary $U\in\mathsf{M}_{m}(\mathbb{C})$ and a diagonal $\Lambda=\lambda I_{p}\oplus\mu I_{m-p}$ such that $A=U\Lambda U^{*}$. Then
$(Y-\lambda X)^{*}(Y-\mu X)$ $=(AX-\lambda X)^{*}(AX-\mu X)$
$=X^{*}(A-\lambda I)^{*}(A-\mu I)X$
$=X^{*}U(\overline{\Lambda}-\overline{\lambda}I)U^{*}U(\Lambda-\mu I)U^{*}X$
$=X^{*}U\begin{bmatrix}0_{p}&0\\ 0&(\overline{\mu}-\overline{\lambda})I_{m-p}\end{bmatrix}\begin{bmatrix}(\lambda-\mu)I_{p}&0\\ 0&0_{m-p}\end{bmatrix}U^{*}X$
[…]
($\Leftarrow$) Suppose that $(Y-\lambda X)^{*}(Y-\mu X)=0$. Let $E=Y-\mu X$ and $F=Y-\lambda X$ and observe that $F^{*}E=0=E^{*}F$, so $\operatorname{col}E$ and $\operatorname{col}F$ are orthogonal subspaces. Moreover,
$X=(\lambda-\mu)^{-1}(E-F)\quad\text{and}\quad Y=(\lambda-\mu)^{-1}(\lambda E-\mu F).$
Let $P=E(E^{*}E)^{\dagger}E^{*}$ and $Q=F(F^{*}F)^{\dagger}F^{*}$, which are, respectively, orthogonal projections onto the respective orthogonal subspaces $\operatorname{col}E$ and $\operatorname{col}F$. If $\operatorname{rank}E+\operatorname{rank}F<n$, let $R$ be the orthogonal projection onto $(\operatorname{col}E\oplus\operatorname{col}F)^{\perp}$; otherwise, let $R=0_{m}$. Let $A=\lambda P+\mu Q+\lambda R$, which is normal with spectrum $\{\lambda,\mu\}$ [3, § 14.9]. Then
$AE$ $=(\lambda P+\mu Q+\lambda R)E=\lambda PE+\mu QE+\lambda RE=\lambda PE=\lambda E,$
$AF$ $=(\lambda P+\mu Q+\lambda R)F=\lambda PF+\mu QF+\lambda RF=\mu QF=\mu F,$
and
$AX=(\lambda-\mu)^{-1}A(E-F)=(\lambda-\mu)^{-1}(AE-AF)=(\lambda-\mu)^{-1}(\lambda E-\mu F)=Y.$
(b) If $X$, $Y$, $\lambda$, and $\mu$ are real, then the construction in part (a) creates real matrices $E$, $F$, $P$, $Q$, and $A$. ∎
Let $X,Y\in\mathsf{M}_{m\times n}(\mathbb{C})$. An orthogonal projection is a normal matrix with spectrum $\{0,1\}$. Theorem 10.1 says that there is an orthogonal projection $A$ such that $AX=Y$ if and only if $(Y-X)^{*}(Y-0X)=Y^{*}Y-X^{*}Y=0$. This is the condition in Theorem 8.1.
A reflection is a normal matrix with spectrum $\{1,-1\}$.Theorem 10.1 says that there is a reflection $A$ such that $AX=Y$ if and only if $C=(Y+X)^{*}(Y-X)=0$. Observe that $C=(Y^{*}Y-X^{*}X)+(X^{*}Y-Y^{*}X)$ is the sum of a Hermitian matrix and a skew-Hermitian matrix, and that $C=0$ if and only if both its Hermitian part and its skew-Hermitian part are zero. This is the condition in Theorem 7.1.
Acknowledgement. We thank Michael A. Dritschel for pointing out that [6, Corollary 1] gives a version of Theorem 5.1 that is valid for bounded linear operators on a Hilbert space. We thank the anonymous referee for detailed comments on the initial draft.